(a) Using the Pythagorean Theorem, we can state that Show that d)d) and (b) We need the potential energy of the marble. This is due entirely to gravity, so U=mgy + C to be such that U-0 when ф 0. where we can choose C to be anything that makes our lives easier. We want What is C? Show that with this choice, 4b (using the expression for s found in (a).)
Answer #1

a) The cycloid's curve is given by

l)(c) + sin(c) ) r

and

y=b (1-\cos (\phi ))

Differentiate x with respect to phi to obtain

dx=b (1+\cos (\phi ))d\phi

Differentiate y with respect y to obtain

dy= b \sin (\phi )\,d\phi

Hence

ds^2=dx^2+dy^2= b^2 (1+\cos (\phi ))^2\,d\phi^2+b^2 \sin ^2(\phi )\,d\phi^2

This can be simplified to

z? b211 +2 cos(4) s2-b(1+cos(p)

Now use the identity

\cos^2\phi+\sin^2\phi=1

to obtain

ds^2=b^2[2+2\cos(\phi)]d\phi^2=2b^2[1+\cos(\phi)]d\phi^2

Now use the identity

1+\cos(\phi)=2\cos^2(\phi/2)

to obtain

ds^2=2b^2[1+\cos(\phi)]d\phi^2=4b^2\cos^2(\phi/2)d\phi^2

Now take the square root on both sides

ds=2b\cos(\phi/2)\,d\phi

and integrate to obtain s

s=2b\int_0^\phi\cos(\phi/2)\,d\phi=4b\sin(\phi/2)

b) The potential energy with C = 0 becomes

U=mgy=b \, m \,g(1-\cos (\phi ))

Now use the identity

(1-\cos (\phi ))=2\sin^2(\phi/2)

to obtain

2b m g sin-(4) /2 16 b2 sin(6/2 8b 8b

Now plug in

s=4b\sin(\phi/2)

and write U in terms of s

U=\frac{mg}{8b}\left(4b\sin(\phi/2) \right )^2=\frac{mg}{2\times 4 b }s^2=\frac{1}{2}\frac{mg}{4b}s^2